# The Power of LLM-Generated Synthetic Data for Stance Detection in Online Political Discussions

### v1

@@ 5 Discussion
Limitations and future work.
One limitation of our approach is the fact that we have to fine-tune a separate model for each question. While this leads to excellent results, a common approach is to fine-tune a single (and thus more general) model for several questions (like Baseline). Having observed that the synthetic data distribution differs (sometimes greatly) for each question, we feel that the extra training is justifiable and necessary. The computing time is not a concern in our case, since fine-tuning the BERT model takes less than a minute even with a synthetic dataset size of $M=1000$ on a reasonable GPU (NVIDIA A100). Future work, could study whether the underlying distribution of the synthetic data can serve as a proxy to combine real unlabelled data from different questions with similar (compatible) underlying distributions. This might enable us to train capable general stance detection models.
Conclusion.
@@
In this work, we presented how to improve stance detection models for online political discussions utilizing LLM-generated synthetic data: (i) we showed that fine-tuning with synthetic data related to the question improves the performance of the stance detection model. After analysing the embeddings of the synthetic data, we attribute this to the LLM-generated synthetic data aligning well with the real data for the given question. (ii) We reduce the labelling effort by selecting the most informative samples informed by synthetic data with SQBC. The synthetic data are separated by construction into the positive and negative classes, which makes finding the most informative samples more straightforward, since we can count the labels of the k-nearest synthetic neighbours to determine classification ambiguity. (iii) Fine-tuning with synthetic data can be further enhanced by adding the most informative samples, outperforming the method that uses all true labels. Concluding, our findings demonstrate the potential of synthetic data to significantly improve the efficiency and effectiveness of stance detection in online political discussions.
Appendix A Broader Impact
Our work has the potential to improve the efficiency and effectiveness of stance detection in online political discussions. This can be beneficial for social media platforms, news outlets, and political campaigns, where the detection of stance can help to understand the sentiment of the users. The use of synthetic data in this case is not critical for the user, since the synthetic data is only used to improve the model’s performance. Stance detection could also be used to for bad purposes, e.g., to manipulate the public opinion. However, our work does not contribute to this, since we only focus on improving the model’s performance, i.e., the model’s ability to detect stance.
Appendix B Visualizations
B.1 Visualizing the synthetic data
@@
We visualize the synthetic data together with the real world data for $M=1000$ and $M=200$ in Figures 5 and 6. We plot the data points of the synthetic data in blue and red for the positive and negative samples, respectively. The means of the real world data are plotted as a regular polygon with 8 sides. We observe that the synthetic data extends the real world data, which we consider a factor as to why fine-tuning with synthetic data is effective in online political discussions. Also, the larger the synthetic dataset size, the more the synthetic data matches the distribution of the real world data since for $M=200$ (see Figure 6) the mean are not as well aligned with the synthetic data. Furthermore, the positive and negative samples are well separated, which we attribute to having pre-trained the BERT-model on $\mathcal{D}_{\text{train}}$ of the X-Stance dataset, giving the prior knowledge about the stance detection task.
@@
B.2 Visualizing the query strategies of the active learning methods
We visualize the selected samples of SQBC, CAL and Random query strategies for $M=1000$ and $M=200$ in Figures 7 and 8. We plot the selected samples of the unlabelled data in green. The positive and negative synthetic data samples are plotted in blue and red, respectively. The selected samples are highlighted in orange. We observe that SQBC selects the unlabelled samples that are mostly in between the two classes of the synthetic data. This is the expected behaviour since we select the samples where the classification score is ambiguous. For Random, the range of selected samples is broad: some similar samples between the two classes like SQBC are selected, but also within class samples that are not covered by the synthetic data set. This explains why random selection works well with a large synthetic dataset, since it further extends the decision boundary of the model. For the smaller synthetic dataset $M=200$, the random selection is not as effective, since the selected samples are spread out over the whole data space and not necessarily in between the two classes as with the larger synthetic dataset. Finally, CAL selects samples similar to SQBC, but mostly tends to select samples from only one class.
@@ B.2 Visualizing the query strategies of the active learning methods
Appendix C Additional Experimental Details
C.1 Compute and Runtime
C.2 Translation of the X-Stance dataset for synthetic data generation
In Figure 9 we show the pipeline for translating the X-Stance dataset for synthetic data generation. We start with a question $q$ from the X-Stance test dataset and translate the question to English with a NLLB-330M model [NLLB Team et al., 2022]. Then we let the Mistral-7B model generate synthetic data, i.e., comments for the translated question. The generated comments are then translated back to German to be used for fine-tuning the model in our experiments.
C.3 Overview of used datasets
Figure 9: Overview of the pipeline for active learning with synthetic data: We start with a question $q$ from the X-Stance test dataset and translate the question to English with a NLLB-330M model [NLLB Team et al., 2022]. Then we let the Mistral-7B model generate synthetic data, i.e., comments for the translated question. The generated comments are then translated back to German to be used for fine-tuning the model in our experiments.
Baseline
Baseline + Synth
Oracle
Oracle + Synth
SQBC
SQBC + Synth
CAL
CAL + Synth
Random ✓(randomly selected)
Random + Synth ✓(randomly selected)
Appendix D Dataset
D.1 Chosen questions and their distribution
We present the $10$ chosen questions for our experiments in Table 2. We show the original questions in German and their corresponding English translations by the translation model. Furthermore, we also show the (60 / 40 ) train/test split for each question in Figure 10. We chose 10 questions that reflect the overall distribution of $\mathcal{D}^{(q)}_{\text{test}}$. We chose unbalanced, balanced and low sample size questions to evaluate the effectiveness of our approach.

### v2

@@ 6 Discussion
Potential impact of this work. An apparent advantage of our approach is the possiblity of divding the synthetic data generation and the stance model fine-tuning. The former can be outsourced to dedicated infrastructure, while with the latter fine-tuning and inference is accessible even for smaller organisations with fewer resources. Considering the large amount of topics in participation processes, the ability to generate synthetic data or to reduce labelling effort with SQBC further increases the benefits for smaller organisations that can’t afford large scale data collection or labelling efforts.
Limitations. One limitation of our approach is that we fine-tune a separate model for each question. While this leads to good results, a common approach is to fine-tune a single (and thus more general) model for several questions (like pre-training Baseline). However, visualising the synthetic data in Figure 3 and Appendix A, we observe that the underlying data distribution differs (sometimes greatly) for each question, which strongly suggests that each question benefits from fine-tuning a different model. This also aligns well with the per topic setting of online (political) discussions, considering that lightweight stance detection models can be fine-tuned in less than a minute even with a synthetic dataset size of $M=1000$ on a reasonable GPU (NVIDIA A100). Another concern are biases that could be potentially introduced through the synthetic data. We addressed this in Section 4 by comparing the distributions of the synthetic data and real world data. We argue that analysing potential biases that could be introduced to the stance detection model through the synthetic data is easier in a single-question setting. In a multiple-question setting data from other topics could introduce biases into the model that are harder to detect. While effective for online discussions, due to the above observations, synthetic data quality should be assessed separately for different use cases.
Future work Despite this work focusing on smaller interpretable models, we believe future work should investigate why the Mistral-7B model performs poorly on the X-Stance dataset. We shared our thoughts around this in Section 5.2, but this requires more extensive study. Another avenue for future work could be about leveraging the synthetic data distribution to generate more novel synthetic data. For instance, with synthetic data as reference we could learn a generative model which enforces a certain (distributional) distance to the synthetic reference distribution.
@@
Conclusion. In this work, we presented how to improve stance detection models for online political discussions utilizing LLM-generated synthetic data: (i) we showed that fine-tuning with synthetic data related to the question improves the performance of the stance detection model. (ii) We attribute this to the LLM-generated synthetic data aligning well with the real data for the given question, while showing that the BERT model requires data that is more content aligned than structured. (iii) Fine-tuning with synthetic data can be further enhanced by adding the most informative samples which are identified by using the synthetic data as reference. This proves more effective than fine-tuning on all true labels, while using considerably less manually labelled samples, thus reducing labelling effort.
Appendix A Visualizations
A.1 Visualizing the synthetic data
@@
We visualize the synthetic data together with the real world data for $M=1000$ and $M=200$ in Figures 6 and 7. We plot the data points of the synthetic data in blue and red for the positive and negative samples, respectively. The means of the real world data are plotted as a regular polygon with 8 sides. We observe that the synthetic data extends the real world data, which we consider a factor as to why fine-tuning with synthetic data is effective in online political discussions. Also, the larger the synthetic dataset size, the more the synthetic data matches the distribution of the real world data since for $M=200$ (see Figure 7) the mean are not as well aligned with the synthetic data. Furthermore, the positive and negative samples are well separated, which we attribute to having pre-trained the BERT-model on $\mathcal{D}_{\text{train}}$ of the X-Stance dataset, giving the prior knowledge about the stance detection task.
@@
A.2 Visualizing the query strategies of the active learning methods
We visualize the selected samples of SQBC, CAL and Random query strategies for $M=1000$ and $M=200$ in Figures 8 and 9. We plot the selected samples of the unlabelled data in green. The positive and negative synthetic data samples are plotted in blue and red, respectively. The selected samples are highlighted in orange. We observe that SQBC selects the unlabelled samples that are mostly in between the two classes of the synthetic data. This is the expected behaviour since we select the samples where the classification score is ambiguous. For Random, the range of selected samples is broad: some similar samples between the two classes like SQBC are selected, but also within class samples that are not covered by the synthetic data set. This explains why random selection works well with a large synthetic dataset, since it further extends the decision boundary of the model. For the smaller synthetic dataset $M=200$, the random selection is not as effective, since the selected samples are spread out over the whole data space and not necessarily in between the two classes as with the larger synthetic dataset. Finally, CAL selects samples similar to SQBC, but mostly tends to select samples from only one class.
@@ A.2 Visualizing the query strategies of the active learning methods
Appendix B Detailed Results
Fine-tuning with synthetic data
M=0 M=200 M=500 M=1000
Baseline + Synth 0.693 0.712 0.718 0.723
True Labels + Synth 0.727 0.745 0.746 0.770
Table 4: Tabular version of Figure 2
Fine-tuning with most informative samples selected with synthetic data
M=200 M=500 M=1000
$10\%$ $25\%$ $50\%$ $75\%$ $10\%$ $25\%$ $50\%$ $75\%$ $10\%$ $25\%$ $50\%$ $75\%$
CAL 0.693 0.697 0.705 0.714 0.692 0.694 0.707 0.718 0.692 0.696 0.708 0.720
Random 0.693 0.696 0.705 0.719 0.693 0.695 0.706 0.715 0.692 0.695 0.706 0.715
SQBC 0.693 0.697 0.709 0.722 0.692 0.700 0.711 0.722 0.692 0.698 0.712 0.721
Table 5: Results of only training with most informative samples.
Fine-tuning with most informative samples and synthetic data
M=200 M=500 M=1000
$10\%$ $25\%$ $50\%$ $75\%$ $10\%$ $25\%$ $50\%$ $75\%$ $10\%$ $25\%$ $50\%$ $75\%$
CAL+Synth 0.713 0.715 0.727 0.732 0.711 0.721 0.732 0.748 0.695 0.715 0.747 0.749
Random+Synth 0.716 0.720 0.724 0.734 0.723 0.735 0.730 0.748 0.724 0.746 0.754 0.756
SQBC+Synth 0.715 0.723 0.731 0.735 0.714 0.726 0.744 0.750 0.721 0.737 0.753 0.747
Table 6: Tabular version of Figure 4
Appendix C Synthetic Data
C.1 Analysing the synthetic data distribution
As described in Section 5.2, we analyse the distribution of the synthetic data in comparison to the real world data. In the following tables we provide samples of the interquartile ranges of the entropy distribution. Tables 7 and 8 contain the same samples in English and German respectively. Table 9 contain comments from the real data distribution.
We see that that the synthetic data samples are longer than the real data samples and contain more reserved language and practically no emotional language. In comparison, the real data samples are shorter and concise with some samples in the lower entropy range containing more emotional language or little information. This however is common for online political discussions. In any case, we see that the underlying notions behind the comments are similar in both the synthetic and real data samples.
Distribution Range Comment
Minimum Sure, I can help you with that! Please write your message and I will assist you with it.
$0\%-25\%$ I strongly disagree with the idea that the residence permit for migrants from non-EU/EFTA countries should be linked to the fulfillment of binding integration agreements throughout Switzerland. While it’s important that migrants integrate into Swiss society and contribute to the country, I believe that it unfair to tie their ability to live and work in Switzerland to their compliance with certain conditions. It could lead to discrimination against some migrants who may struggle to meet these requirements or face cultural barriers. Instead, I propose a more nuanced approach that takes into account the unique circumstances and needs of each individual migrant, while still promoting integration and equality for all.
$25\%-50\%$ As someone who strongly believes in the importance of immigration policies that promote successful integration and contribution to society, I fully support the idea of linking residence permits for migrants from non-EU/EFTA countries to the fulfillment of binding integration agreements throughout Switzerland. This approach ensures that migrants committed to making Switzerland their home and contributing positively to the country’s well-being can enjoy full legal recognition and protection while being held accountable for fulfilling their responsibilities. By creating an incentive system where residency and benefits are directly tied to compliance with integration agreements, we can foster a culture of shared responsibility and mutual benefit between incoming migrants and Swiss society as a whole.
$50\%-75\%$ Sure, I can help you generate a response for this scenario. From your perspective, you might write something like: I strongly believe that the residence permit for migrants from non-EU/EFTA countries should be linked to the fulfillment of binding integration agreements throughout Switzerland. As someone who values the diversity and richness that migration brings to our country, I think it’s important that we have clear expectations and requirements for those coming here to live and work. Integration agreements should include things like language proficiency, cultural awareness, and contributions to community life. By linking these agreements to the residency permit, we can ensure that migrants are fully committed to integrating into Swiss society and contributing to its well-being. Of course, there may be some concerns or challenges with this approach, but I believe that the benefits far outweigh any potential drawbacks. A strong and diverse immigration policy is crucial for the future of Switzerland, and by setting high standards for integration, we can attract the best and brightest from around the world.
$75\%-100\%$ I believe that requiring residence permits for migrants from non-EU/EFTA countries to be linked to the fulfillment of binding integration agreements throughout Switzerland is a fair and necessary policy. By doing so, we can ensure that migrants who come to Switzerland have a clear understanding of their responsibilities and the requirements they need to meet in order to live and work here. This also helps to promote a sense of community and integration among all residents of Switzerland, regardless of their place of birth or nationality. Additionally, it allows us to carefully screen and select immigrants who are likely to be successful and positive contributing members of our society. Overall, I believe that this policy will benefit both migrants and Swiss citizens alike by promoting mutual respect and understanding.
Maximum I would argue that linking the residence permit for migrants from non-EU/EFTA countries to the fulfillment of mandatory integration agreements across Switzerland is too burdensome and unreasonable. This creates an unfair situation in which people who come to Switzerland to work or study are essentially forced to make a long-term commitment before they even have access to the legal rights they need to live and work here comfortably.The Commission has asked the Commission to review the conditions set out in Regulation (EC) No. 1225/2009 for the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No. 1225/2009 of the Council concerning the application of Regulation (EC) No.I believe that the government should instead focus on creating more flexible and streamlined processes to facilitate the integration of migrants into Swiss society without imposing unnecessary restrictions or requirements.
Table 7: English comments sampled from the interquartire ranges of the entropy distribution of the synthetic data.
Distribution Range Comment
Minimum Schreiben Sie Ihre Nachricht, ich helfe Ihnen.
$0\%-25\%$ Ich bin mit der Vorstellung, dass die Aufenthaltserlaubnis für Migranten aus Nicht-EU/EFTA-Ländern an die Erfüllung verbindlicher Integrationsvereinbarungen in der gesamten Schweiz gebunden werden sollte, voll und ganz nicht einverstanden. Es ist zwar wichtig, dass Migranten in die Schweizer Gesellschaft integriert werden und einen Beitrag für das Land leisten, aber ich glaube, es ist unfair, ihre Fähigkeit, in der Schweiz zu leben und zu arbeiten, an die Einhaltung bestimmter Bedingungen zu knüpfen. Es könnte zu Diskriminierung einiger Migranten führen, die möglicherweise Schwierigkeiten haben, diese Anforderungen zu erfüllen oder kulturellen Hindernissen ausgesetzt sind. Ich schlage stattdessen einen differenzierteren Ansatz vor, der die einzigartigen Umstände und Bedürfnisse jedes einzelnen Migranten berücksichtigt und gleichzeitig Integration und Gleichberechtigung für alle Migranten fördert.
$25\%-50\%$ Als jemand, der fest an die Bedeutung von Einwanderungspolitiken glaubt, die eine erfolgreiche Integration und einen Beitrag zur Gesellschaft fördern, unterstütze ich die Idee, Aufenthaltsgenehmigungen für Migranten aus Nicht-EU/EFTA-Ländern mit der Erfüllung verbindlicher Integrationsvereinbarungen in der gesamten Schweiz zu verknüpfen. Dieser Ansatz gewährleistet, dass Migranten, die sich verpflichtet haben, die Schweiz zu ihrer Heimat zu machen und einen positiven Beitrag zum Wohlergehen des Landes leisten, vollständige rechtliche Anerkennung und Schutz genießen können, während sie gleichzeitig für die Erfüllung ihrer Pflichten zur Rechenschaft gezogen werden. Durch die Schaffung eines Anreizsystems, bei dem Aufenthalt und Leistungen unmittelbar an die Einhaltung der Integrationsvereinbarungen gebunden sind, können wir eine Kultur der gemeinsamen Verantwortung und des gegenseitigen Nutzens zwischen den einwandernden Migranten und der Schweizer Gesellschaft als Ganzes fördern.
$50\%-75\%$ Sicher, ich kann Ihnen helfen, eine Antwort für dieses Szenario zu generieren Aus Ihrer Sicht könnten Sie etwa schreiben: Ich bin der festen Überzeugung, dass die Aufenthaltserlaubnis für Migranten aus Nicht-EU/EFTA-Ländern mit der Erfüllung verbindlicher Integrationsvereinbarungen in der gesamten Schweiz verbunden sein sollte Als jemand, der die Vielfalt und den Reichtum schätzt, die Migration unserem Land bringt, denke ich, dass es wichtig ist, dass wir klare Erwartungen und Anforderungen an diejenigen haben, die hierher kommen, um zu leben und zu arbeiten. Die Integrationsvereinbarungen sollten Aspekte wie Sprachkenntnisse, kulturelles Bewusstsein und Beiträge zum Gemeinschaftsleben umfassen Durch die Verknüpfung dieser Vereinbarungen mit der Aufenthaltserlaubnis können wir sicherstellen, dass Migranten sich voll und ganz der Integration in die Schweizer Gesellschaft und dem Wohlergehen der Schweizer Gesellschaft verschrieben haben. Natürlich kann dieser Ansatz einige Bedenken oder Herausforderungen mit sich bringen, aber ich glaube, dass die Vorteile weit überwiegen alle möglichen Nachteile Eine starke und vielfältige Einwanderungspolitik ist für die Zukunft der Schweiz von entscheidender Bedeutung. Durch die Festlegung hoher Integrationsstandards können wir die Besten und Klügsten aus der ganzen Welt anziehen.
$75\%-100\%$ Ich glaube, dass die Anerkennung der Aufenthaltsgenehmigung für Migranten aus Nicht-EU/EFTA-Ländern an die Erfüllung verbindlicher Integrationsvereinbarungen in der gesamten Schweiz geknüpft ist, eine faire und notwendige Politik Auf diese Weise können wir sicherstellen, dass Migranten, die in die Schweiz kommen, ein klares Verständnis ihrer Verantwortlichkeiten und der Anforderungen haben, die sie erfüllen müssen, um hier zu leben und zu arbeiten. Dies trägt auch dazu bei, ein Gefühl der Gemeinschaft und Integration unter allen Einwohnern der Schweiz zu fördern, unabhängig von ihrem Geburtsort oder ihrer Staatsangehörigkeit. Darüber hinaus können wir so Migranten sorgfältig auswählen, die wahrscheinlich erfolgreiche und positive Mitglieder unserer Gesellschaft sind. Ich glaube, dass diese Politik sowohl den Migranten als auch den Schweizer Bürgern zugute kommen wird, indem sie gegenseitigen Respekt und Verständnis fördert
Maximum Ich würde argumentieren, dass die Verknüpfung der Aufenthaltserlaubnis für Migranten aus Nicht-EU/EFTA-Ländern mit der Erfüllung verbindlicher Integrationsvereinbarungen in der gesamten Schweiz zu belastend und unvernünftig ist Es entsteht eine ungerechte Situation, in der Menschen, die zur Arbeit oder zum Studium in die Schweiz kommen, im Wesentlichen gezwungen sind, eine langfristige Verpflichtung einzugehen, bevor sie überhaupt Zugang zu den gesetzlichen Rechten haben, die sie benötigen, um hier bequem zu leben und zu arbeiten. Die Kommission hat die Kommission aufgefordert, die in der Verordnung (EG) Nr. 1225/2009 festgelegten Bedingungen für die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. 1225/2009 des Rates über die Anwendung der Verordnung (EG) Nr. Ich glaube, dass die Regierung sich stattdessen auf die Schaffung flexiblerer und schlankerer Prozesse konzentrieren sollte, um die Integration von Migranten in die Schweizer Gesellschaft zu erleichtern, ohne unnötige Einschränkungen oder Anforderungen aufzuerlegen.
Table 8: Corresponding German comments sampled from the interquartire ranges of the entropy distribution of the synthetic data.
Distribution Range Comment
Minimum muss die integrationsvereinbarungen noch studieren
$0\%-25\%$ Wer sich anpasst und korrekt verhaltet darf auch hier bleiben.
$25\%-50\%$ Integrationsvereinbarungen nicht generell verhängen, sondern individuell anordnen, wenn die Person auch nach längerer Anwesenheit in der Schweiz Integrationsdefizite hat oder bei neu zugezogenen Personen, wenn mit Anpassungsschwierigkeiten zu rechnen ist. Keine Standardformulare, keine Automatismen.
$50\%-75\%$ Ich zweifle an der Umsetzbarkeit solcher Methoden, eines "Vertrags". Es ist anstrebenswert, dass ein guter Integrationsablauf statt finden kann. Allerdings ist dafür immer eine Gesellschaft notwendig, die Personen aufnehmen und integrieren will, wie auch Personen die diese Möglichkeit wahrnehmen können und wollen.
$75\%-100\%$ Ich finde die Idee Integrationsmassnahmen verbindlich zu machen sinnvoll. Sprachkurse sind ein ideales Instrument dafür. Es darf aber nicht sein, dass gewisse Forderungen zur Voraussetzung werden für die Aufenthaltserlaubnis. Nicht jeder Mensch lernt gleich schnell eine Sprache oder kann sich gleich schnell integrieren. Missbräuche und Diskriminierungen können auftreten. Dies muss verhindert werden. Ich wäre eher für einen obligatorischen Kostenlosen-Sprachkurs.
Maximum Heikler Punkt, im Prinzip JA für Alle, nur - auch gleiche Rechte/Pflichten für alle - warum sollen die EU-EFTA-Bürger keine solchen Integrationsvereinbarungen abschliessen. Der bildungsferne Rumäne fährt besser als der hochausgebildete Chinese/US-Amerikaner/IT-Spezialist aus Indien. Die ganze Einwanderung muss im Sinne der MEI umgesetzt werden.
Table 9: Corresponding German comments sampled from the interquartire ranges of the entropy distribution of the real data.
C.2 Translated Data Samples
We show the translated questions used for synthetic data generation in Table 10 and some samples of generated comments in 12. We see that the questions are translated correctly and synthetic data can be generated for both favor and against stances.